OpenStudy (anonymous):

Can someone really skilled in number theory help out? I am not talking about a school homework, i am trying to solve a problem for fun. Also if you can give links I will just read. I googled but couldn't find anything interesting.

OpenStudy (amoodarya):

if a,b,c are prime numbers and a^2-b^2=c find a,b,c

OpenStudy (anonymous):

I guess 3, 2 and 5 are prime numbers? This is just by trial and error though :)

OpenStudy (amoodarya):

ok good but it has a solution look (a+b)(a-b) =c*1 all of theme are prime so a+b=c and a-b=1 because a-b <a+b (in N) just two prime number can be a,b because a=b+1 only 2 is even prime number but all the others ore odd numbers so a=3 ,b=2 hence c=2+3

OpenStudy (anonymous):

i am talking about a completely different stuff here ppl. Have you heard of the sigma function? Sum of the divisors of a number?

OpenStudy (anonymous):

please don't post random stuff here

OpenStudy (anonymous):

No, not familiar with that, sorry.

OpenStudy (amoodarya):

now try other problem how many prime number like "p" exist such that 5p+1 is sqr of natural number

OpenStudy (amoodarya):

how many prime in sequence of a(n)=n!+3

OpenStudy (anonymous):

I can't see an original question by you mivanov.

OpenStudy (anonymous):

It goes like example for 12 we have divisors 1,2,3,4,6,12 and we use sigma2 which is the sum of these squared 1^2+2^2+3^2+4^2+6^2+12^2. so we call that sigma2 and then we want to calculate sum of sigma2 for n ranging from 1 to 10^15 and modulo 10^9 the result

OpenStudy (anonymous):

it's this problem. i made a program that calculates it to like 1 000 000 https://projecteuler.net/problem=401 but it's just slow for numbers above 1 000 000

OpenStudy (anonymous):

amoodarya can you help me with problem? give me some hints about the SIGMA2 function please

OpenStudy (anonymous):

@amoodarya

OpenStudy (amoodarya):

can you list 10^15 divisors ? if yes from 2^4 *5^4 sima2 func modulo 10^9 is zero so you have to compute 1^2+2^2 +5^2 +10^2 +..10^4 mod 10^9

OpenStudy (amoodarya):

do you know the formula that give sum of divisors?

OpenStudy (amoodarya):